A long straight copper wire of radius R carries a free current I. The current is distributed uniformly over the cross-section of the wire. Find the magnetizing field H outside and inside the wire


1
Expert's answer
2022-01-24T01:14:07-0500

"H_1=0,"

"H_2=\frac B{\mu_0}=\frac{\mu I}{2\pi R}."